OpenStudy (anonymous):

Select the correct solution set. 50 ≥ 15x {x | x ≥ 10/3} {x | x ≤ 10/3} {x | x > 10/3}

OpenStudy (anonymous):

@ninjaslice

OpenStudy (ninjaslice):

I would take it, that it would be A, but I may be wrong, I doubt i'm wrong though lol.

OpenStudy (anonymous):

They are equal so it would be A.

OpenStudy (anonymous):

if you are greater than your brother then your brother is smaller than you.

OpenStudy (ninjaslice):

@Alina_Dubrovnik Thanks for verifying my answer!

OpenStudy (anonymous):

\[if~ 50 \ge 15x,then~15x \le 50\]

OpenStudy (anonymous):

\[x \le ?\]